OpenStudy (anonymous):

A newspaper boy was delivering the papers to three of his clients. He traveled 2x miles from his home to the first client, and 2 miles to the second client. From the second client's house to the next client is 3x miles. All in all, he traveled 13 miles from his house to the third client's house. How far is the first client's house from his house amd the third client's house to the second client's house?

OpenStudy (anonymous):

2x = home to 1st house 2 = 1st to 2nd 3x = 2nd to 3rd total = 13mi so you can get this equation: 2x+2+3x=13 solve for x 5x=11 x=11/5 so from house to 1st client = 2(11/5) = 22/5miles and 2nd to 3rd client = 3(11/5)=33/5